OpenStudy (anonymous):

Evaluate 7 + 52 • 4.

OpenStudy (anonymous):

214 35 128 107

OpenStudy (anonymous):

its not 128

OpenStudy (anonymous):

anf its not 52 is 5^2

OpenStudy (anonymous):

Firstly find the value of \(5^2\) ?? Tell me what is value of 5*5 = ??

OpenStudy (anonymous):

Has your teacher ever taught you PEMDAS? It's an acronym.

OpenStudy (anonymous):

Firstly solve for exponent then multiply the result you got and then add 7 to it..

OpenStudy (anonymous):

\[\implies 7 + (5 \times 5) \times 4 \implies 7 + (25) \times 4 \implies \color{green}{7 + 100} = ??\]

OpenStudy (anonymous):

If you still have any questions, just ask. I'm sure one of us could walk you through how to do any sort of problem like this, step by step.

OpenStudy (anonymous):

okay jackoff i know how to do pemdas i had a problem with this one question if your gunna be rude on help me fluttering quire

OpenStudy (anonymous):

I'm sorry, I didn't mean to come across as rude. If you know how to do these kinds of problems, that's great!